1 + cos x/sin x = cot x/2
can you please help me

Answer by heraklea(1)

1+(cosx)/(sinx)=ctg(x/2)
1+cos((2*(x/2))/sin((2*(x/2))=(cos(x/2))/((sin(x/2))
1+(((cos(x/2))*2-((sin(x/2))*2)/(2*sin(x/2)*cos(x/2))=(cos(x/2))/(sin(x/2))
x=pi/2